OpenStudy (anonymous):

give vector v = 1/2. Is this a unit vector? if not, make into one

OpenStudy (amistre64):

um, the i and j seem out of place there; those tend to be reserved for a more equation looking vector notation

OpenStudy (anonymous):

ok try this then...u=1/2(sqrt(3)i-j)

OpenStudy (amistre64):

thats better :)

OpenStudy (amistre64):

so, what we have is: \[\vec v=<\frac{\sqrt{3}}{2},-\frac{1}{2}>\] which looks to me like points on the unit circle

OpenStudy (amistre64):

sqrt((3+1)/4) = 1

OpenStudy (anonymous):

ok.... so is it a unit vector as is?

OpenStudy (amistre64):

as is, its a unit vector; since its magnitude = 1

OpenStudy (anonymous):

thanks quick question if you could. When we say directional derivative, that means DuF right.

OpenStudy (amistre64):

http://mathworld.wolfram.com/DirectionalDerivative.html in a sense I would say yes

OpenStudy (anonymous):

What i mean is, when we say find the dirctional derrivative its reffering to that.